OpenStudy (anonymous):

consider rolling two dice.what is the probability that the sum is odd given that one dice shows a 3?

OpenStudy (julie):

well, you could brute force this

OpenStudy (julie):

so assuming it's a six sided dice...

OpenStudy (julie):

there are six possibilities for the other die

OpenStudy (julie):

3+1 = 4 [even] 3+2 = 5 [odd] 3+ 3 = 6 [even] 3+ 4 = 7 [odd] 3 + 5 = 8 [even] 3 + 6 = 9 [odd]

OpenStudy (julie):

so you know the roll of one of the dice is a three

OpenStudy (julie):

looks to me there are 3 options that get you odd, 3 that give you even

OpenStudy (julie):

in sum

OpenStudy (julie):

(so 3/6 possibilities for the second die to make an odd number) = 50%

OpenStudy (anonymous):

ohhhhh ok i get it

OpenStudy (julie):

that's generally how these work - you want to consider <success case>/<total possibilities> - especially given all possibilities are equally likely e.g. a fair die

OpenStudy (anonymous):

oh wow i wish my teacher would have put it that way

OpenStudy (julie):

good to hear, good luck!

OpenStudy (anonymous):

thanks goin